# Joint Task Offloading and Routing in Wireless Multi-hop Networks Using Biased Backpressure Algorithm

### v1

@@ Abstract
A significant challenge for computation offloading in wireless multi-hop networks is the complex interaction among traffic flows in the presence of interference. Existing approaches often ignore these key effects and/or rely on outdated queueing and channel state information. To fill these gaps, we reformulate joint offloading and routing as a routing problem on an extended graph with physical and virtual links. We adopt the state-of-the-art shortest path-biased Backpressure routing algorithm, which allows the destination and the route of a job to be dynamically adjusted at every time step based on network-wide long-term information and real-time states of local neighborhoods. In large networks, our approach achieves smaller makespan than existing approaches, such as separated Backpressure offloading and joint offloading and routing based on linear programming.
@@ 1 Introduction
Computation offloading often focuses on reducing the makespan (job initialization to completion time) and/or energy consumption of tasks by deciding where and when each job should be processed as well as how to upload the job description to the server [1]. A popular approach is to use mixed-integer programming (MIP) [20, 21, 22, 23, 24, 25, 26], which seeks to optimally schedule individual jobs, subject to the constraints of their dependency relationships and the residual capacities of servers and communication channels. MIP has been applied to scenarios with complex task dependency [25, 23] and where servers are only one hop away from clients [25, 26, 24]. In multi-hop offloading, however, MIP faces the challenges of 1) NP-hardness in finding a solution and 2) timeliness in collecting the residual capacities of servers multiple hops away from the scheduler. The mean-field linear programming (LP) relaxation [21, 22, 23], despite more tractable, often overlooks the key effects of queueing and link scheduling for managing the contentions among traffic flows and transceivers [27].
When jobs are recurrent and can be independently processed by a server, stochastic network optimization (SNO) is a better option [28, 29, 30, 31, 32, 33, 34], in which decisions are driven by the queueing state information (QSI) of clients, servers, and relays. Compared to MIP, SNO can handle the network stochasticity by queueing the jobs, while satisfying the mean-field constraints (queue stability) by Lyapunov optimization. Originated from routing, SNO is suitable for joint offloading and routing [30, 31, 32, 33, 34]. However, since the destination of a job is determined at the decision phase, these schemes still suffer from the changes in QSI between the time it was collected for offloading decision and the moment the job arrives to its destination.
@@
1) For the first time, we incorporate the idea of virtual sink in the field of network flow [35] to enable joint offloading and routing in multi-hop networks under a unified routing framework. In particular, we model the computing at a server as sending jobs to a virtual sink over a virtual link – an operation compatible with communications.
@@
Fig. 1: Graph modeling: (a) Task offloading in a wireless multi-hop network where a client may have two types of tasks. (b) Extended graph and per-commodity queues for the example in (a). (c) Job scheduling for two types of jobs with equal priority at a computing node $i$ with four processors, the computing capacity of the server are modeled as two virtual links, with link rate equal to the corresponding service rate.
@@ 2 System Model and Problem Formulation
The conflict graph, ${\mathcal{G}}^{c}=({\mathcal{E}},{\mathcal{H}})$, describes the conflict relationship between links (here, we re-use the notation ${\mathcal{E}}$ referring its elements as vertices) and is defined as follows: each vertex $e\in{\mathcal{E}}$ corresponds to a link in ${\mathcal{G}}^{n}$ and each undirected edge $(e_{1},e_{2})\in{\mathcal{H}}$ indicates a conflict between links $e_{1},e_{2}\in{\mathcal{E}}$ in ${\mathcal{G}}^{n}$. Two links could be in conflict due to either 1) interface conflict, i.e., two links share a wireless device with only one wireless transceiver; or 2) wireless interference, i.e., their incident devices are within a certain distance such that their simultaneous transmission will cause the outage probability to exceed a prescribed level. In this paper, we assume $\mathcal{G}^{c}$ to be known, possibly by each node monitoring the wireless channel [18], or through more sophisticated approaches [36].
@@
In joint offloading and routing, we make two types of decisions: 1) where each job should be processed, e.g., at its source client (locally) or a specific remote server; and 2) through which route the job should be uploaded to the specified server. Our objective is to minimize the average makespan (communication time and computing time) of all tasks across the network. In the mean field, we can model each task $c_{m}$ as a fluid flow with a rate of $\lambda_{m}^{c}$ injected to the network via client $m$, and formulate the problem as optimizing the static non-negative flow rate assignments $\{f_{m}^{c}(e)|e\in{\mathcal{E}},m\in{\mathcal{M}},c\in{\mathcal{C}}\}$, or $\{f_{m}^{c}(e)\}$ in short, to minimize the total makespan of all tasks:
@@
where (1c) is for all $m\in{\mathcal{M}},c\in{\mathcal{C}},v\in{\mathcal{V}}$. In (1), $u(e)$ and $u^{c}(v)$ are respectively the unit costs of sending flow over link $e$ and processing job flow of type $c$ via node $v$, $\psi(v)$ and $\psi(e)$ are respectively the long-term computing and communication capacities of node $v$ and link $e$ under given scheduling policies, ${\mathcal{N}}(v)$ is the set of immediate neighbors of node $v$ on ${\mathcal{G}}^{n}$, $g_{m}^{c}(v)$ is the processing rate of flow $c_{m}$ on node $v$, and $h^{c}(v)$ is the unit computing capacity demanded by job flow of type $c$ on $v$. The constraints in (1) are explained as follows: (1b) defines the average makespan of task $c_{m}$ as the sum of communication and computing costs across all the links and nodes under policy $\{f_{m}^{c}(e)\}$. (1c) states the flow conservation law that the inflow of any node must equal to its outflow, and flows must be only on the existing links $e\in{\mathcal{E}}$. (1d) states that the total processing rate of flow $c_{m}$ in the network must equal to its arrival rate, i.e., all jobs must be processed. (1e) and (1f) respectively state that computing on node $v$ and communication over link $e$ are limited by their long-term capacities. (1g) limits all processing rates, unit capacity demands, and unit costs to be non-negative real values.
@@ 4 Numerical experiments
Our joint offloading and routing (joint_SP-BP) and other benchmark schemes are evaluated on 10 randomly generated wireless ad-hoc networks with 100 nodes, as exemplified in Fig. 3. Each connectivity graph ${\mathcal{G}}^{n}$ has a 3-tier structure: first, draw $k\in\{3,4,5,6\}$ with uniform distribution, and create a clique of $k$ core servers (tier-1); next, $4k$ tier-2 nodes ($80\%$ edge servers and $20\%$ dedicated relays) in $k$ groups are added, nodes in the same group are connected to the same core server, and connected with each other with a probability of $0.7$; lastly, each of the $100-5k$ clients (tier-3) is connected to one (or two with a probability of $0.1$) random server(s) (an edge server with a probability of $10/(41k)$ or a core server with a probability of $1/(41k)$). The conflict graph ${\mathcal{G}}^{c}$ is the line graph of ${\mathcal{G}}^{n}$. The base service rates $\bm{\mu}$ are generated as follows: for servers, $|{\mathcal{S}}|$ random values are drawn from a Pareto distribution with a shape of 2 and a scale of 8, and the $s$th largest element is assigned to the server added to the network in the $s$th order; for a client $m$, $\bm{\mu}_{m}\in\mathbbm{U}(8,12)$.
@@
and jobs are taken to their destinations by SP-BP routing. The server backlog $Q_{i}^{(c)}(t)$ is assumed to be instantly available to client $m$ at $t$, and the biases for offloading and routing are the same ones in (3). 2) BP_SP-BP [28, 29, 30]: same as SPBP_SP-BP but the biases in (8) (offloading only) are zeros. 3) joint_LP: the probabilities for a job of task $c_{m}$ being sent to a neighbor or processed at its residing node are proportional to the static optimal rates $\{f^{c}_{m}(e)\}^{*}\cup\{g^{c}_{m}(v)\}^{*}$, which are found by solving the LP formulation in (1), assuming that $\psi(v)=\bm{\mu}_{v}^{0}$, $\psi(e)={\mathbf{r}}_{e}/d^{c}(e)$, $u(v)=1/\psi(v),u(e)=1/\psi(e)$, where $d^{c}(e)$ is the degree of link $e$ on the conflict graph ${\mathcal{G}}^{c}$.
We first compare joint_SP-BP, BP_SP-BP, and SPBP_SP-BP under the setting described before. In Figs. 2(a) and 2(b), we plot the makespan and number of hops a job traveled before completion for each of the three schemes. BP_SP-BP performs poorly by only relying on QSI, as evidenced by the large makespan and hop distance of jobs. SPBP_SP-BP enhances BP_SP-BP by considering the distances ($B_{m}^{(i)}$) to and computing capacities ($B_{i}^{(c)}$) of servers in offloading decisions, substantially reducing the offloading distance and makespan. Our joint_SP-BP further outperforms SPBP_SP-BP, confirming the benefit of allowing the route of each job to the virtual sink (e.g., where the job should be processed) to be dynamically adjusted based on accurate network state, rather than deciding it before routing. The small makespan under joint_SP-BP shows that it alleviates the last packet problem by merging bursty tasks into only two commodities rather than $|{\mathcal{S}}|$ in the separated schemes.
Next, we present the makespan under joint_SP-BP, SPBP_SP-BP, and joint_LP in Fig. 2(c), for a modified setting with only one task type ${\mathcal{C}}=\{0\}$, service rates $\bm{\mu}^{0}=0.125\bm{\mu}$, all tasks are streaming, and everything else the same as the previous test. The joint_LP performs similarly to joint_SP-BP for low traffic load $l=1.0$, but as $l$ increases, it degrades quickly and underperforms SPBP_SP-BP, and could not find a valid solution for $l\geq 2.0$. This is due to the underlying assumptions of joint_LP, which deviate further from reality as traffic load increases, and that the static probabilistic policy from joint_LP cannot react to real-time network congestion as ours.
@@ 5 Conclusions
By modeling the computing at a node as sending jobs to virtual sinks over virtual links and leveraging the throughput optimality of Backpressure routing, our fully distributed approach can simultaneously achieve offloading, load balancing, and routing, closing the information gaps in conventional approaches. Moreover, our approach avoids the excessive communication overhead, the dependency on a centralized scheduler, and unrealistic assumptions in mean-field LP approaches. Potential future directions include incorporating downlink flows, task priorities, and complex task workflows.

### v3

@@ Abstract
A significant challenge for computation offloading in wireless multi-hop networks is the complex interactions among traffic flows in the presence of interference. Existing approaches often ignore these key effects and/or rely on outdated queueing and channel state information. To fill these gaps, we reformulate joint offloading and routing as a routing problem on an extended graph with physical and virtual links. We adopt the state-of-the-art shortest path-biased Backpressure routing algorithm, which allows the destination and the route of a job to be dynamically adjusted at every time step based on network-wide long-term information and real-time states of local neighborhoods. In large networks, our approach achieves smaller makespan than existing approaches, such as separated Backpressure offloading, and joint offloading and routing based on linear programming.
@@ 1 Introduction
Computation offloading often focuses on reducing the makespan (job initialization to completion time) and/or energy consumption of tasks by deciding where and when each job should be processed as well as how to upload the job description to the server [1]. A popular approach is to use mixed-integer programming (MIP) [20, 21, 22, 23, 24, 25, 26], which seeks to optimally schedule individual jobs, subject to the constraints of their dependency relationships and the residual capacities of servers and communication channels. MIP has been applied to scenarios with complex task dependencies [25, 23] and where servers are only one hop away from clients [25, 26, 24]. In multi-hop offloading, however, MIP faces the challenges of 1) NP-hardness in finding a solution and 2) timeliness in collecting the residual capacities of servers multiple hops away from the scheduler. The mean-field linear programming (LP) relaxation [21, 22, 23], is more tractable, but often overlooks the key effects of queuing and link scheduling for managing the contentions among traffic flows and transceivers [27].
When jobs are recurrent and can be independently processed by a server, stochastic network optimization (SNO) is a better option [28, 29, 30, 31, 32, 33, 34], in which decisions are driven by the queuing state information (QSI) of clients, servers, and relays. Compared to MIP, SNO can handle network stochasticity by queuing jobs, while satisfying the mean-field constraints (queue stability) by Lyapunov optimization. Originating from routing, SNO is suitable for joint offloading and routing [30, 31, 32, 33, 34]. However, since the destination of a job is fixed during routing in all the existing schemes [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], shifts in network state (including QSI) may render the initial offloading decision suboptimal by the time the job reaches its destination.
@@
1) For the first time, we incorporate the idea of virtual sink in the field of network flow [35] to enable joint offloading and routing in multi-hop networks under a unified routing framework. In particular, we model computation at a server as sending jobs to a virtual sink over a virtual link – an operation compatible with communications.
@@
Fig. 1: Graph modeling: (a) Task offloading in a wireless multi-hop network where a client may have two types of tasks. (b) Extended graph and per-commodity queues for the example in (a). (c) Job scheduling for two types of jobs with equal priority at a computing node $i$ with four processors, the computing capacity of a server is modeled as two virtual links, with link rate equal to the corresponding service rate.
@@ 2 System Model and Problem Formulation
The conflict graph, ${\mathcal{G}}^{c}=({\mathcal{E}},{\mathcal{H}})$, describes the conflict relationship between links (here, we re-use the notation ${\mathcal{E}}$ referring to its elements as vertices) and is defined as follows: each vertex $e\in{\mathcal{E}}$ corresponds to a link in ${\mathcal{G}}^{n}$ and each undirected edge $(e_{1},e_{2})\in{\mathcal{H}}$ indicates a conflict between links $e_{1},e_{2}\in{\mathcal{E}}$ in ${\mathcal{G}}^{n}$. Two links could be in conflict due to either 1) interface conflict, i.e., two links share a wireless device with only one wireless transceiver; or 2) wireless interference, i.e., their incident devices are within a certain distance such that their simultaneous transmission will cause the outage probability to exceed a prescribed level. In this paper, we assume $\mathcal{G}^{c}$ to be known, possibly by each node monitoring the wireless channel [18], or through more sophisticated approaches [36].
@@
In joint offloading and routing, we make two types of decisions: 1) where each job should be processed, e.g., at its source client (locally) or a specific remote server; and 2) through which route the job should be uploaded to the specified server. Our objective is to minimize the average makespan (communication time and computing time) of all tasks across the network. In the mean field, we can model each task $c_{m}$ as a fluid flow with a rate of $\lambda_{m}^{c}$ injected to the network via client $m$, and formulate the problem as optimizing the static non-negative flow rate assignments $\{f_{m}^{c}(e)|e\in{\mathcal{E}},m\in{\mathcal{M}},c\in{\mathcal{C}}\}$, or simply as $\{f_{m}^{c}(e)\}$, to minimize the total makespan of all tasks:
@@
where (1c) is for all $m\in{\mathcal{M}},c\in{\mathcal{C}},v\in{\mathcal{V}}$. In (1), $u(e)$ and $u^{c}(v)$ are respectively the unit costs of sending flow over link $e$ and processing job flow of type $c$ via node $v$, $\psi(v)$ and $\psi(e)$ are respectively the long-term computing and communication capacities of node $v$ and link $e$ under given scheduling policies, ${\mathcal{N}}(v)$ is the set of immediate neighbors of node $v$ on ${\mathcal{G}}^{n}$, $g_{m}^{c}(v)$ is the processing rate of flow $c_{m}$ on node $v$, and $h^{c}(v)$ is the unit computing capacity demanded by job flow of type $c$ on $v$. The constraints in (1) are explained as follows: (1b) defines the average makespan of task $c_{m}$ as the sum of communication and computing costs across all the links and nodes under policy $\{f_{m}^{c}(e)\}$. (1c) states the flow conservation law that the inflow of any node must equal its outflow, and flows must be only on the existing links $e\in{\mathcal{E}}$. (1d) states that the total processing rate of flow $c_{m}$ in the network must be equal to its arrival rate, i.e., all jobs must be processed. (1e) and (1f) respectively state that computing on node $v$ and communication over link $e$ are limited by their long-term capacities. (1g) limits all processing rates, unit capacity demands, and unit costs to be non-negative real values.
@@ 4 Numerical experiments
Our joint offloading and routing (joint_SP-BP) and other benchmark schemes are evaluated on 10 randomly generated wireless ad-hoc networks with 100 nodes, as exemplified in Fig. 3. Each connectivity graph ${\mathcal{G}}^{n}$ has a 3-tier structure: first, draw $k\in\{3,4,5,6\}$ with uniform distribution, and create a clique of $k$ core servers (tier-1); next, $4k$ tier-2 nodes ($80\%$ edge servers and $20\%$ dedicated relays) in $k$ groups are added, nodes in the $k$th group are all directly connected to the $k$th core server, and connected with each other with a probability of $0.7$; lastly, each of the $100-5k$ clients (tier-3) is connected to one (or two with a probability of $0.1$) random server(s) (an edge server with a probability of $10/(41k)$ or a core server with a probability of $1/(41k)$). The conflict graph ${\mathcal{G}}^{c}$ is the line graph of ${\mathcal{G}}^{n}$. The base service rates $\bm{\mu}$ are generated as follows: for servers, $|{\mathcal{S}}|$ random values are drawn from a Pareto distribution with a shape of 2 and a scale of 8, and the $s$th largest element is assigned to the server added to the network in the $s$th order; for a client $m$, $\bm{\mu}_{m}\in\mathbbm{U}(8,12)$.
@@
and jobs are taken to their destinations by SP-BP routing. The server backlog $Q_{i}^{(c)}(t)$ is assumed to be instantly available to client $m$ at $t$, and the biases for offloading and routing are the ones in (3). 2) BP_SP-BP [28, 29, 30]: same as SPBP_SP-BP but the biases in (8) (offloading only) are zeros. 3) joint_LP: the probabilities for a job of task $c_{m}$ being sent to a neighbor or processed at its residing node are proportional to the static optimal rates $\{f^{c}_{m}(e)\}^{*}\cup\{g^{c}_{m}(v)\}^{*}$, which are found by solving the LP formulation in (1), assuming that $\psi(v)=\bm{\mu}_{v}^{0}$, $\psi(e)={\mathbf{r}}_{e}/d^{c}(e)$, $u(v)=1/\psi(v),u(e)=1/\psi(e)$, where $d^{c}(e)$ is the degree of link $e$ on the conflict graph ${\mathcal{G}}^{c}$.
We first compare joint_SP-BP, BP_SP-BP, and SPBP_SP-BP under the setting described before. In Figs. 2(a) and 2(b), we plot the makespan and number of hops a job travels before completion for each of the three schemes. BP_SP-BP performs poorly by only relying on QSI, as evidenced by the large makespan and hop distance of jobs. SPBP_SP-BP enhances BP_SP-BP by considering the distances ($B_{m}^{(i)}$) to and computing capacities ($B_{i}^{(c)}$) of servers in offloading decisions, substantially reducing the offloading distance and makespan. Our joint_SP-BP further outperforms SPBP_SP-BP, confirming the benefit of allowing the route of each job to the virtual sink (e.g., where the job should be processed) to be dynamically adjusted based on accurate network state, rather than deciding it before routing. The small makespan under joint_SP-BP shows that it alleviates the last packet problem by merging bursty tasks into only two commodities rather than $|{\mathcal{S}}|$ in the separated schemes.
Next, we present the makespan under joint_SP-BP, SPBP_SP-BP, and joint_LP in Fig. 2(c), for a modified setting with only one task type ${\mathcal{C}}=\{0\}$, service rates $\bm{\mu}^{0}=0.125\bm{\mu}$, all tasks are streaming, and everything else the same as the previous test. The joint_LP performs similarly to joint_SP-BP for low traffic load $l=1.0$, but as $l$ increases, it degrades quickly and underperforms SPBP_SP-BP, and could not find a valid solution for $l\geq 2.0$. This is due to the underlying assumptions of joint_LP, which deviate further from reality as traffic load increases, and that the static probabilistic policy from joint_LP cannot react to real-time network congestion as ours does.
@@ 5 Conclusions
By modeling the computing at a node as sending jobs to virtual sinks over virtual links and leveraging the throughput optimality of Backpressure routing, our fully distributed approach can simultaneously achieve offloading, load balancing, and routing, closing the information gaps in conventional approaches. Moreover, our approach avoids excessive communication overhead, dependency on a centralized scheduler, and unrealistic assumptions in mean-field LP approaches. Potential future directions include incorporating downlink flows, task priorities, and complex task workflows.